OpenStudy (rebeccapink6398):

@bruhhh A magician is doing a rope trick with a rope that is 7/8 feet long. During the trick he cuts a piece that is 1/6 of the origianal. How long is the cut piece?...

OpenStudy (anonymous):

Sorry, stepped away! It wants you to multiply the two fractions \[\frac{ 7 }{ 8 }\times \frac{ 1 }{ 6 }=\frac{ 7 }{ 48 }\] Hope that helps ^.^

OpenStudy (rebeccapink6398):

ohhh thank youuu

OpenStudy (anonymous):

lol you're welcome :)